# Investigations of MWISP Clumps: 13CO Clump Source Catalogs and Physical Properties

### v1

@@ 4.2 Distance Assignment of Clumps
The hierarchical distance determination process assigns classifications to all SRs in the FGQ. Due to some SRs having large spatial scales, it is necessary to calculate individual distances for each clump within these SRs. We propagate these SR classifications to individual clumps and assign corresponding distances to each clump. For improved precision, clumps associated with masers and their neighbors receive the corresponding maser distance instead. The distance catalog is detailed in Table 4 of Appendix B.
@@
Table 1 summarizes the KDA solution statistics. A total of 1,718 clumps, approximately 2.4%, have their distances determined using the maser distance, with the KDA method being Matching Maser or located in the SGQ and TGQ. KDA resolution is applied to 72.44% (51,908) of the total, with 73.7% at near-distance and 22.7% at far-distance, in agreement with previous studies (e.g. Eden et al., 2012; Urquhart et al., 2018). Furthermore, 3.6% of the clumps have an unknown KDA classification, with these clumps assigned a tangent distance as indicated in Figure 4. The final catalog of physical parameters encompasses 70,184 clumps, achieving a completeness of 97.94%. Each clump has a validated distance greater than 0.1 kpc, with all unphysical solutions derived from rotation curve extrapolations excluded.
@@
To assess the validation of the distance, we match MWISP clumps with ATLASGAL clumps (Urquhart et al. 2018; defined as ATLASGAL clumps located within MWISP clump masks). This cross-matching yields 2,865 ATLASGAL clumps associated with 1,938 MWISP clumps, of which 1,161 MWISP clumps utilize maser associations for distance classification and 692 employ direct maser distance measurements. As shown in Figure 7, kernel density estimation (KDE) is applied to smoothly estimate the probability density function. Linear regression performed on clumps within regions exceeding 90% KDE values produces a fitted slope of 1.04. Despite some clumps showing anti-correlated distances (a common phenomenon between catalogs due to different distance assignment; Mège et al. 2021), the overall distribution and quantitative analysis reveal strong agreement between both catalogs, validating the SR-centric automated KDA methodology. This study achieves automated KDA solutions, incorporates more maser distances, and enables a maser to determine multiple clumps through the SR, substantially enhancing the reliability of distance classification.
@@
Figure 8 shows the face-on spatial distribution of molecular clumps in the Galactic plane using a Galactocentric coordinate system. Clump number density peaks in the solar neighborhood and declines with increasing heliocentric distance, mainly due to observational bias as discussed in Section 5.1: lower-mass clumps are harder to detect at greater distances. The increased discreteness in longitude ranges $l<15^{\circ}$ and $165^{\circ}<l<195^{\circ}$ stems from larger kinematic distance uncertainties in these regions (Wenger et al., 2018). Within the visualization, a scatter-point arc demarcates the solar circle, while nearby clumps lacking definitive distance classifications have been assigned tangent distances. Our samples exhibit a cavity structure within 3 kpc of the Galactic Center, support previous observational findings of molecular gas deficiency in this region (e.g. Burton, 1976). The molecular gas deficiency in this region may be related to the combined effects of intense nuclear winds impacting the Galactic gas disk (Su et al., 2022) and the dynamical influence of the Galactic bar on gas motion (Su et al., 2024, 2025).
@@
Step Method Number (Percent) Near Far Unknown
@@
(1) Matching Masers 13166 (18.4%) 88.9% 11.1% 0
(2) Outside Sular Circle 5136 (7.2%) 0 100% 0
@@
(8) SGQ and TGQ 19757 (27.56%)
Summary Clumps with KDA 51908 (72.44%) 73.7% 22.7% 3.6%
@@
Figure 9: Relationships between distance-dependent parameters and heliocentric distance. Panels show: (a) MdM-d; (b) $R_{eq}-d$; (c) $\alpha_{\text{vir}}-d$; and (d) $n(\mathrm{H}_{2})-d$. Scatter points from the MWISP project are color-coded according to their KDE values, as indicated by the color bar, reflecting the local data density. MWISP S1 represents a subset of the MWISP survey that covers the similar galactic sector as CHIMPS. Contour lines, labeled for MWISP (cyan), MWISP S1 (pink), CHIMPS (limegreen), SEDIGISM (orange), highlight regions with KDE values exceeding 90%. The red lines represent the reference lines showing the theoretical parameter-distance relationships. For each survey’s high-density region, $\alpha_{0.9}$ denotes the fitted power-law slop of the relationship, presented alongside its uncertainty and $P$.
Figure 10: Probability distribution function histograms for four physical parameters. Panels show: (a) $M$; (b) $R_{eq}$; (c) $\sigma_{v}$; (d) $\alpha_{\text{vir}}$; (e) $n(\mathrm{H}_{2})$; and (f) $\Sigma$. Each sub - plot compares data from MWISP, CHIMPS, and SEDIGISM, with their respective median values marked.
@@ 5 Physical Properties and Scaling Relations
Utilizing multiple transitions of 12CO and 13CO spectral lines, we conduct radiative transfer analysis under the LTE assumption, with detailed calculation methods presented in Appendix C. The analytical process begins with calculating the excitation temperature ($T_{\text{ex}}$) from the brightness temperature ratio of isotopic lines, followed by constraining the optical depth ($\tau$) through radiative transfer balance equations, and subsequently deriving column density ($N(^{13}\mathrm{CO})$) via integration of population distribution over mask velocity space. From $N(^{13}\mathrm{CO})$ measurements, we derive distance-dependent physical parameters: mass ($M$), equivalent radius ($R_{\rm eq}$), surface density ($\Sigma$), and volume density ($n(\mathrm{H}_{2})$). Additionally, we calculate the virial parameter ($\alpha_{\rm vir}$) with velocity dispersion ($\sigma_{v}$) obtained from spectral line measurements to assess the dynamical state of clumps.
The catalog of physical parameters is detailed in Table 5 of Appendix B, with representative parameter maps of exemplary clumps displayed in Appendix E. A quantitative assessment of parameter biases stemming from sensitivity thresholds is presented in Appendix D.1. To minimize measurement uncertainties, we exclude samples with kinematic distances and their corresponding SRs located within the longitude ranges $l<15^{\circ}$ and $165^{\circ}<l<195^{\circ}$. This refined dataset comprises 61,202 clumps, representing 85.4% of the total observed population.
To establish broader context for MWISP survey findings, we compare our results with two other Galactic molecular cloud surveys. The CHIMPS survey (Rigby et al., 2016) mapping 13CO (3–2) and C18O (3–2) emission within $27.8^{\circ}<|l|<46.2^{\circ}$ and $|b|<0.5^{\circ}$, with a spatial resolution of 15′′ and a channel width of 0.5 km s-1, cataloging 4999 13CO (3–2) clumps using FellWalker (Rigby et al., 2019; Berry, 2015); and the SEDIGISM survey (Schuller et al., 2017) observing 13CO (2–1) emission across $60^{\circ}<|l|<+18^{\circ}$ and $|b|<0.5^{\circ}$, with a spatial resolution of 28′′ and a channel width of 0.25 km s-1, cataloging 10663 molecular clouds using SCIMES (Duarte-Cabral et al., 2021; Colombo et al., 2015).
@@ 5.1 Distance-Dependent Observational Biases
Figure 9 evaluates the power-law relationships between distance-dependent physical parameters ($M$, $R_{\text{eq}}$, $\alpha_{\text{vir}}$, and $n_{\text{H}_{2}}$) and heliocentric distance. MWISP S1 represents a subset of the MWISP survey that encompasses the similar galactic sector to CHIMPS, spanning the range $27^{\circ}<l<47^{\circ}$ and $|b|<0.5^{\circ}$. The scatter points are color-coded according to their KDE values, reflecting local data density, while contour lines highlight high-density regions with KDE values exceeding 90%. The fitted power-law slopes ($\alpha_{0.9}$) are calculated through orthogonal distance regression (Rigby et al., 2019; Virtanen et al., 2020; Feng et al., 2024), and Pearson correlation coefficients ($P$) quantify the correlation significance. The red lines represent theoretical distance-parameter relationships as described in Appendix C.2 for comparison.
Figures 9(a) and (b) reveal parallel behavior in mass and size parameters across different surveys. Figure 9(a) shows positive correlations between mass and distance with slopes steeper than the theoretical geometric effects $M\propto d^{2}$ expectation (MWISP: $\alpha_{0.9}=2.781\pm 0.009$, $P=0.83$; MWISP S1: $\alpha_{0.9}=5.046\pm 0.079$; CHIMPS: $\alpha_{0.9}=5.514$; SEDIGISM: $\alpha_{0.9}=4.12$). Figure 9(b) demonstrates that MWISP exhibits excellent agreement with the theoretical $R_{\text{eq}}\propto d$ relationship ($\alpha_{0.9}=1.023\pm 0.002$, $P=0.91$), showing minimal deviation from expected scaling, while MWISP S1, CHIMPS, and SEDIGISM display departures ($\alpha_{0.9}=1.197\pm 0.01$, $1.817$ and $1.516$, respectively). These deviations arise from observational selection effects that systematically bias measurements toward larger values at greater distances: (1) sensitivity-limited Malmquist bias that excludes faint, low-mass clumps (Roman-Duval et al., 2010); (2) beam dilution and surface brightness thresholds that preferentially detect bright, extended sources; and (3) angular resolution limitations causing beam confusion where multiple small clumps appear as single massive sources.
@@
These correlations between clump parameters and distance (e.g. Roman-Duval et al., 2010; Duarte-Cabral et al., 2021; Rani et al., 2023) are primarily driven by observational constraints rather than intrinsic physical variations. The universal nature of these biases reflects limitations in sensitivity, angular resolution, and surface brightness detection thresholds that affect millimeter and submillimeter surveys. MWISP S1 shows steeper slopes than the complete MWISP sample across all parameters, indicating that observational selection effects vary across different galactic regions. The comparison between MWISP S1 and CHIMPS, which probe the similar galactic regions and eliminate potential confusion from different galactic environments, reveals that MWISP S1 exhibits more moderate slopes than CHIMPS.
MWISP clumps exhibit slopes closer to theoretical geometric effects reference lines across most parameters, potentially indicating less severe additional distance-related effects than other catalogs. Beyond differences in instrumentation and observational coverage, these reduced biases likely originate from both distance calculation methods and the FacetClumps detection algorithm, which adopts pixel distance-independent parameter constraints and extracts homogeneous clump characteristics, yielding a more uniform population of molecular structures across varying distances. The resulting dataset provides a valuable resource for investigating the intrinsic physical properties of clump structures.
@@ 5.2 Clump Physical Parameter Distributions
Figure 10 presents probability distribution function (PDF) histograms for six physical parameters across different molecular cloud/clump catalogs, revealing distinct populations of structures that trace different physical regimes.
Figure 10(a) shows that SEDIGISM detects the most massive structures (median $1136.37M_{\odot}$), while CHIMPS ($183.61M_{\odot}$) and MWISP ($225.78~M_{\odot}$) trace less massive objects. Figure 10(b) differentiates structural information, with SEDIGISM identifying the most spatially extended objects (median $2.18$ pc), followed by MWISP ($1.72$ pc) and CHIMPS ($1.23$ pc). SEDIGISM’s catalog with the SCIMES algorithm specifically targets cloud-scale structures, accounting for its higher median $M$ and $R_{\text{eq}}$, while both MWISP and CHIMPS detect smaller clumps within larger complexes.
@@
These distinct parameter distributions illustrate how the catalogs systematically target different components of the molecular clump population: MWISP characterizes extended, moderately massive structures with intermediate virial parameters, low velocity dispersions, and low densities; CHIMPS captures similar-sized regions with comparable masses but higher virial parameters and intermediate densities; SEDIGISM detects the most massive cloud-scale structures with larger extents and higher densities. The magnitude of these differences exceeds what would be expected from distance-dependent biases alone, highlighting the genuine physical distinctions between structures targeted by each catalog.
Figure 11: Scaling relations between physical parameters of different surveys. Panels show: (a) $M$-$R_{\rm eq}$; (b) $n({\rm H_{2}})$-$R_{\rm eq}$; (c) $\alpha_{\rm vir}$-$R_{\rm eq}$; (d) $\sigma_{v}$-$R_{\rm eq}$; (e) $\sigma_{v}$-$M$; and (f) $\sigma_{v}$-$R_{\rm eq}\Sigma$. The color scale represents the KDE of MWISP clumps. Colored dashed lines indicate power-law fits to the main distributions outlined by corresponding contours, with derived power-law indices ($\alpha_{0.9}\pm$ uncertainty) and $P$ labeled for each survey.

### v2

@@ 4.2 Distance Assignment of Clumps
The hierarchical distance determination process assigns classifications to all SRs in the FGQ. Due to some SRs having large spatial scales, it is necessary to calculate individual distances for each clump within these SRs. We propagate these SR classifications to individual clumps and assign corresponding distances to each clump. For improved precision, clumps associated with masers and their neighbors receive the corresponding maser distance instead. The distance catalog is detailed in Table B.2 of Appendix B.
@@
Table 1 summarizes the KDA solution statistics. A total of 1718 clumps, approximately 2.4%, have their distances determined using the maser distance, with the KDA method being Matching Maser or located in the SGQ and TGQ. KDA resolution is applied to 72.44% (51,908) of the total, with 73.7% at near-distance and 22.7% at far-distance, in agreement with previous studies (e.g. Eden et al., 2012; Urquhart et al., 2018). Furthermore, 3.6% of the clumps have an unknown KDA classification, with these clumps assigned a tangent distance as indicated in Figure 4. The final catalog of physical parameters encompasses 70,184 clumps, achieving a completeness of 97.94%. Each clump has a validated distance greater than 0.1 kpc, with all unphysical solutions derived from rotation curve extrapolations excluded.
@@
To assess the validation of the distance, we match MWISP clumps with ATLASGAL clumps (Urquhart et al. 2018; defined as ATLASGAL clumps located within MWISP clump masks). This cross-matching yields 2865 ATLASGAL clumps associated with 1938 MWISP clumps, of which 1161 MWISP clumps utilize maser associations for distance classification and 692 employ direct maser distance measurements. As shown in Figure 7, kernel density estimation (KDE) is applied to smoothly estimate the probability density function. Linear regression performed on clumps within regions exceeding 90% KDE values produces a fitted slope of 1.04. Despite some clumps showing anti-correlated distances (a common phenomenon between catalogs due to different distance assignments; Mège et al. 2021), the overall distribution and quantitative analysis reveal strong agreement between both catalogs, validating the SR-centric automated KDA methodology. This study achieves automated KDA solutions, incorporates more maser distances, and enables a maser to determine multiple clumps through the SR, substantially enhancing the reliability of distance classification.
@@
Figure 8 shows the face-on spatial distribution of molecular clumps in the Galactic plane using a Galactocentric coordinate system. Clump number density peaks in the solar neighborhood and declines with increasing heliocentric distance, mainly due to observational bias as discussed in Section 5.1: lower-mass clumps are harder to detect at greater distances. The increased discreteness in longitude ranges $l<15^{\circ}$ and $165^{\circ}<l<195^{\circ}$ stems from larger kinematic distance uncertainties in these regions (Wenger et al., 2018). Within the visualization, a scatter-point arc demarcates the solar circle, while nearby clumps lacking definitive distance classifications have been assigned tangent distances. Our samples exhibit a cavity structure within 3 kpc of the Galactic center, supporting previous observational findings of molecular gas deficiency in this region (e.g., Burton, 1976). The molecular gas deficiency in this region may be related to the combined effects of intense nuclear winds impacting the Galactic gas disk (Su et al., 2022) and the dynamical influence of the Galactic bar on gas motion (Su et al., 2024, 2025).
@@
Step Method Number (Percentage) Near Far Unknown
@@
(1) Matching Masers 13,166 (18.4%) 88.9% 11.1% 0
(2) Outside Solar Circle 5136 (7.2%) 0 100% 0
@@
(8) SGQ and TGQ 19,757 (27.56%)
Summary Clumps with KDA 51,908 (72.44%) 73.7% 22.7% 3.6%
@@
Figure 9: Relationships between distance-dependent parameters and heliocentric distance. Panels show (a) MdM-d; (b) $R_{eq}-d$; (c) $\alpha_{\text{vir}}-d$; and (d) $n(\mathrm{H}_{2})-d$. Scatter points from the MWISP project are color-coded according to their KDE values, as indicated by the color bar, reflecting the local data density. MWISP S1 represents a subset of the MWISP survey that covers a similar Galactic sector to CHIMPS. Contour lines, labeled for MWISP (cyan), MWISP S1 (pink), CHIMPS (lime green), SEDIGISM (orange), highlight regions with KDE values exceeding 90%. The red lines represent the reference lines showing the theoretical parameter-distance relationships. For each survey’s high-density region, $\alpha_{0.9}$ denotes the fitted power-law slope of the relationship, presented alongside its uncertainty and $P$.
Figure 10: Probability distribution function histograms for six physical parameters. Panels show (a) $M$; (b) $R_{eq}$; (c) $\sigma_{v}$; (d) $\alpha_{\text{vir}}$; (e) $\Sigma$; and (f) $n(\mathrm{H}_{2})$. Each subplot compares data from MWISP, CHIMPS, and SEDIGISM, with their respective median values marked.
@@ 5 Physical Properties and Scaling Relations
Utilizing multiple transitions of 12CO and 13CO spectral lines, we conduct radiative transfer analysis under the LTE assumption, with the detailed calculation methods presented in Appendix C. The analytical process begins with calculating the excitation temperature ($T_{\text{ex}}$) from the brightness temperature ratio of isotopic lines, followed by constraining the optical depth ($\tau$) through radiative transfer balance equations, and subsequently deriving the column density ($N(^{13}\mathrm{CO})$) via integration of the population distribution over the mask velocity space. From the $N(^{13}\mathrm{CO})$ measurements, we derive distance-dependent physical parameters: mass ($M$), equivalent radius ($R_{\rm eq}$), surface density ($\Sigma$), and volume density ($n(\mathrm{H}_{2})$). Additionally, we calculate the virial parameter ($\alpha_{\rm vir}$) with velocity dispersion ($\sigma_{v}$) obtained from spectral line measurements to assess the dynamical state of clumps.
The catalog of physical parameters is detailed in Table B.3 of Appendix B, with representative parameter maps of exemplary clumps displayed in Appendix E. A quantitative assessment of parameter biases stemming from sensitivity thresholds is presented in Appendix D.1. To minimize measurement uncertainties, we exclude samples with kinematic distances and their corresponding SRs located within the longitude ranges $l<15^{\circ}$ and $165^{\circ}<l<195^{\circ}$. This refined dataset comprises 61,202 clumps, representing 85.4% of the total observed population.
To establish broader context for MWISP survey findings, we compare our results with those of two other Galactic molecular cloud surveys. The CHIMPS survey (Rigby et al., 2016) mapped 13CO (3–2) and C18O (3–2) emission within $27.8^{\circ}<|l|<46.2^{\circ}$ and $|b|<0.5^{\circ}$, with a spatial resolution of 15′′ and a channel width of 0.5 km s-1, and cataloged 4999 13CO (3–2) clumps using FellWalker (Rigby et al., 2019; Berry, 2015); and the SEDIGISM survey (Schuller et al., 2017) observing 13CO (2–1) emission across $60^{\circ}<|l|<+18^{\circ}$ and $|b|<0.5^{\circ}$, with a spatial resolution of 28′′ and a channel width of 0.25 km s-1, and cataloged 10,663 molecular clouds using SCIMES (Duarte-Cabral et al., 2021; Colombo et al., 2015).
@@ 5.1 Distance-Dependent Observational Biases
Figure 9 evaluates the power-law relationships between distance-dependent physical parameters ($M$, $R_{\text{eq}}$, $\alpha_{\text{vir}}$, and $n_{\text{H}_{2}}$) and heliocentric distance. MWISP S1 represents a subset of the MWISP survey that encompasses a similar Galactic sector to CHIMPS, spanning the range $27^{\circ}<l<47^{\circ}$ and $|b|<0.5^{\circ}$. The scatter points are color-coded according to their KDE values, reflecting local data density, while contour lines highlight high-density regions with KDE values exceeding 90%. The fitted power-law slopes ($\alpha_{0.9}$) are calculated through orthogonal distance regression (Rigby et al., 2019; Virtanen et al., 2020; Feng et al., 2024), and Pearson correlation coefficients ($P$) quantify the correlation significance. The red lines represent theoretical distance-parameter relationships as described in Appendix C.2 for comparison.
Figures 9(a) and (b) reveal parallel behavior in mass and size parameters across different surveys. Figure 9(a) shows positive correlations between mass and distance with slopes steeper than the theoretical geometric effect $M\propto d^{2}$ expectation (MWISP: $\alpha_{0.9}=2.781\pm 0.009$, $P=0.83$; MWISP S1: $\alpha_{0.9}=5.046\pm 0.079$; CHIMPS: $\alpha_{0.9}=5.514$; SEDIGISM: $\alpha_{0.9}=4.12$). Figure 9(b) demonstrates that MWISP exhibits excellent agreement with the theoretical $R_{\text{eq}}\propto d$ relationship ($\alpha_{0.9}=1.023\pm 0.002$, $P=0.91$), showing minimal deviation from expected scaling, while MWISP S1, CHIMPS, and SEDIGISM display departures ($\alpha_{0.9}=1.197\pm 0.01$, $1.817$, and $1.516$, respectively). These deviations arise from observational selection effects that systematically bias measurements toward larger values at greater distances: (1) sensitivity-limited Malmquist bias that excludes faint, low-mass clumps (Roman-Duval et al., 2010); (2) beam dilution and surface brightness thresholds that preferentially detect bright, extended sources; and (3) angular resolution limitations causing beam confusion where multiple small clumps appear as single massive sources.
@@
These correlations between clump parameters and distance (e.g., Roman-Duval et al., 2010; Duarte-Cabral et al., 2021; Rani et al., 2023) are primarily driven by observational constraints rather than intrinsic physical variations. The universal nature of these biases reflects limitations in sensitivity, angular resolution, and surface brightness detection thresholds that affect millimeter and submillimeter surveys. MWISP S1 shows steeper slopes than the complete MWISP sample across all parameters, indicating that observational selection effects vary across different Galactic regions. The comparison between MWISP S1 and CHIMPS, which probe the similar Galactic regions and eliminate potential confusion from different Galactic environments, reveals that MWISP S1 exhibits more moderate slopes than CHIMPS.
MWISP clumps exhibit slopes closer to theoretical geometric effect reference lines across most parameters, potentially indicating less severe additional distance-related effects than other catalogs. Beyond differences in instrumentation and observational coverage, these reduced biases likely originate from both distance calculation methods and the FacetClumps detection algorithm, which adopts pixel distance-independent parameter constraints and extracts homogeneous clump characteristics, yielding a more uniform population of molecular structures across varying distances. The resulting dataset provides a valuable resource for investigating the intrinsic physical properties of clump structures.
@@ 5.2 Clump Physical Parameter Distributions
Figure 10 presents probability distribution function histograms for six physical parameters across different molecular cloud/clump catalogs, revealing distinct populations of structures that trace different physical regimes.
Figure 10(a) shows that SEDIGISM detects the most massive structures (median $1136.37~M_{\odot}$), while CHIMPS ($183.61~M_{\odot}$) and MWISP ($225.78~M_{\odot}$) trace less massive objects. Figure 10(b) differentiates structural information, with SEDIGISM identifying the most spatially extended objects (median $2.18$ pc), followed by MWISP ($1.72$ pc) and CHIMPS ($1.23$ pc). SEDIGISM’s catalog with the SCIMES algorithm specifically targets cloud-scale structures, accounting for its higher median $M$ and $R_{\text{eq}}$, while both MWISP and CHIMPS detect smaller clumps within larger complexes.
@@
These distinct parameter distributions illustrate how the catalogs systematically target different components of the molecular clump population: MWISP characterizes extended, moderately massive structures with intermediate virial parameters, low velocity dispersions, and low densities; CHIMPS captures similar-sized regions with comparable masses but higher virial parameters and intermediate densities; and SEDIGISM detects the most massive cloud-scale structures with larger extents and higher densities. The magnitude of these differences exceeds what would be expected from distance-dependent biases alone, highlighting the genuine physical distinctions between structures targeted by each catalog.
Figure 11: Scaling relations between physical parameters of different surveys. Panels show (a) $M$-$R_{\rm eq}$; (b) $n({\rm H_{2}})$-$R_{\rm eq}$; (c) $\alpha_{\rm vir}$-$R_{\rm eq}$; (d) $\sigma_{v}$-$R_{\rm eq}$; (e) $\sigma_{v}$-$M$; and (f) $\sigma_{v}$-$R_{\rm eq}\Sigma$. The color scale represents the KDE of MWISP clumps. Colored dashed lines indicate power-law fits to the main distributions outlined by corresponding contours, with derived power-law indices ($\alpha_{0.9}\pm$ uncertainty) and $P$ labeled for each survey.